PLEASE HELP

A snow maker at a resort pumps 190 kg of lake water per minute and sprays it into the air above a ski run. The water droplets freeze in the air and fall to the ground, forming a layer of snow. If all of the water pumped into the air turns to snow, and the snow cools to the ambient air temperature of -5.5
Answer #1

change in temp of water from 14.9C to 0C

Q1 = mcdT = 190*4186*14.9 = 1.185x10^7 J

freezing of water at 0C

Q2 = mLf = 190*334000 = 6.346x10^7 J

change in temp of water from 0C to -5.5C

Q3 = mcdT = 190*2000*5.5 = 0.209x19^7 J

Q = Q1 + Q2 + Q3

Q = 1.185x10^7 + 6.346x10^7 + 0.209x19^7

Q = 7.74x10^7 J

Heat released per minute = 7.74x10^7/60 = 1.29x10^6